OpenStudy (anonymous):

what is negative square root of 25/81

OpenStudy (unklerhaukus):

\[-\sqrt\frac{25}{81}\]\[=-\frac{\sqrt{25}}{\sqrt{81}}\]\[=-\frac{\sqrt{5^2}}{\sqrt{9^2}}\]\[=\]

OpenStudy (anonymous):

thanks!